OpenStudy (anonymous):

Please help Determine the delivery radius for your shop. Draw a point on a coordinate plane where your shop will be located. Create two different radii lengths from your shop, and construct the circles that represent each delivery area. How much area will each delivery radius cover? Write the equation for each circle created.

OpenStudy (anonymous):

judging by this question, you don't have any restrictions, so just draw two circles in a xy-coordinate plane, which you can easily translate into the equation of a circle. \[f(x,y)=(x-u)^2+(y-v)^2=r^2 \] where (u,v) are the x and y respective coordinates of the center of your circle and r is the radius.